Steam at 15 bar is passed through a throttle valve with a pressure drop of 12 bar. The temperature after throttling is 150 °C.

Q.4 The dryness fraction before throttling is


1
Expert's answer
2021-04-12T01:47:16-0400

Let ,

p1= inlet pressure=15 ×10-6 Pa

x1 = Quality of steam at inlet

T1= inlet temperature

h1 = Inlet specific enthalpy

p2= Exit pressure=3 ×10-6 Pa

T2= Temperature at exit=150 C

h2=specific Enthalpy at exit.

At T2= 150 C and P2=3 bar we can find specific Enthalpy from steam table,

h2=2571 KJ/kg

Since, during throttling process, enthalpy remains constant,

h1=h2=2571 kJ/kg

At inlet pressure P1=15 bar , hf=844.56 kJ/kg, hfg= 1946.4 kJ/kg

h1=hf+X hfg=2571

844.56+X (844.56)=2571.4

X=0.887

Answer:

Dryness fraction of steam at inlet

X=0.887